OpenStudy (anonymous):

Solve the inequality \[\frac{4}{x^2+4}<0\] and express the solution as an interval.

OpenStudy (anonymous):

numerator is always positive, and so is the denominator, so no solution

OpenStudy (anonymous):

Yeah I knew there is no way possible this can be less than 0....

OpenStudy (anonymous):

\(4>0\) and \(x^2\geq 0\) for all \(x\) and so \(x^2+4\geq 4\) for all \(x\) and therefore \(\frac{4}{x^2+4}>0\) for all x

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

wait, how'd you get that last part?

OpenStudy (anonymous):

what last part?

OpenStudy (anonymous):

how is \[x^2 \ge 0\]

OpenStudy (anonymous):

because it is a square

OpenStudy (anonymous):

if you square a positive number, the answer is positive if you square a negative number, the answer is also positive the only way to get something not positive is to square 0 that is why \(x^2\geq0\)

OpenStudy (anonymous):

Oh... okay, wasn't reading it properly... thanks once again :)